# Effective models for dense vortex lattices in the Kitaev honeycomb model

### 22 Dec 2023 v1

@@ I Introduction
The characterization of free fermion phases via topological invariants represents a major achievement in the understanding of quantum matter [1, 2, 3, 4]. Given the spatial dimension, these invariants may be determined through an analysis of global symmetries such as time-reversal (TRS), particle-hole, and chiral symmetries, in a classification scheme known as the tenfold way. In 2006, Kitaev introduced an exactly solvable spin-1/2 model on a honeycomb lattice that can be mapped to a free Majorana-fermion problem in a static $\mathbb{Z}_{2}$ gauge field [5]. The gauge field results in the possibility of static $\pi$ fluxes, called vortices, in each plaquette and the problem separates into independent vortex sectors. In the isotropic limit and in the vortex-free sector where the ground state lies (see below for details), the spectrum is gapless, but acquires a gap and a nonzero Chern number $\nu$ when TRS is broken [5]. To break TRS while preserving the integrability of the model, Kitaev introduced a three-spin term that results in a next-nearest neighbor hopping of the Majorana fermions. The corresponding model, closely related to Haldane’s model [6], has $\nu=\pm 1$ and has been the subject of many studies (see, e.g., Refs. [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]).
@@
The “duals” to these triangular vortex configurations correspond to dense vortex configurations, with density $\rho=(n-1)/n$, where the plaquettes that do not contain vortices form a triangular pattern against a background of plaquettes that all contain vortices. These dense vortex configurations have recently been studied numerically in Refs. [14, 15], and can be divided into two families according to their behavior with respect to translation and inversion symmetry. We shall refer to these two families as “dual even” (if $n$ is even such as $\rho=3/4,11/12,15/16,...$) and “dual odd” (if $n$ is odd such as $\rho=2/3,8/9,11/12,...$) in the following [see Fig. 1(b) and (c)] [8, 15]. The dual even family exhibits gapped phases with even Chern numbers, separated by extended gapless phases (spin metals with a Majorana-Fermi surface [23, 24]), whereas the dual odd family has gapped phases with even or odd Chern numbers, separated by gapless points. These dual configurations are not describable by the effective model introduced in Refs. [12, 13].
In this article, we construct two different effective models describing the dual even and dual odd families. These models describe the hopping of pairs of coupled Majorana fermions on a triangular lattice, where the lattice sites correspond to the absences of vortices within a vortex sea. These effective models allow us to understand the variety of Chern numbers and the nature of gapless phases observed. The article is organized as follows: in Sec. II we briefly introduce the Kitaev honeycomb model in the Majorana fermion language and discuss the behaviors of the vortex-free and vortex-full sectors. In Sec. III we identify three distinct symmetry families for direct and dual vortex configurations based on translation and inversion symmetries. In Sec. IV we show that when the low-energy bands of the Kitaev honeycomb model are well-separated from the high-energy bands, they may be accurately described by an effective tight-binding model for all three of these families. We first review the construction of the low-energy model discussed in Refs. [12, 13] for the direct case, and then introduce new effective models for the dual even and dual odd cases. To demonstrate the accuracy of these models, we directly compare their integrated density of states (IDOS) with that of the original Kitaev honeycomb model. Finally, in Sec. V, we discuss our results, conclude and give perspectives. Appendices provide details on specific points such as the validity of the effective models (App. A) and some possible dispersion relations of vortex bands (App. B).
@@ II Kitaev honeycomb model
Here the $u_{jk}$ parameters are $\mathbb{Z}_{2}$ gauge variables defined such that $u_{jk}=-u_{kj}=\pm 1$ for each link of the lattice. We assume here that the coupling $J$ is the same in all three directions, corresponding to the isotropic limit of the original spin model. This Hamiltonian describes noninteracting Majorana fermions coupled to a static $\mathbb{Z}_{2}$ gauge field. The term $J$ describes the nearest-neighbor hopping, whereas $\kappa$ denotes the strength of the three-spin term that breaks TRS, corresponding to next-nearest neighbor hopping (see Fig. 2). In the following, without loss of generality, we will assume that $J$ and $\kappa$ are nonnegative.
For each plaquette $p$ of the honeycomb lattice, we define the $\mathbb{Z}_{2}$ variable $w_{p}=\prod_{(j,k)\in p}u_{jk}={\rm e}^{i\phi_{p}}$, where $\phi_{p}$ is the flux threading the plaquette. Here, for each plaquette, the pair of sites $(j,k)$ is taken such that $k$ always belongs to one sublattice of the honeycomb lattice, and $j$ belongs to the other. A value of $w_{p}=-1$ ($+1$) indicates a flux $\phi_{p}=\pi$ (00) in the plaquette $p$, meaning the presence (absence) of a vortex. Two sets of $\mathbb{Z}_{2}$ link variables $u_{jk}$ defined on the lattice are said to be equivalent if they correspond to the same configuration of plaquette variables, or “vortex configuration.” In analyzing Eq. (1), one can then restrict to a particular vortex sector labelled by the pattern of $w_{p}$’s. In Fig. 1 we show several examples of vortex configurations, where the $w_{p}=-1$ plaquettes are shaded grey.
Direct vortex configurations [see Fig. 1 (a)] are defined by two vectors $\bm{A}_{1}$ and $\bm{A}_{2}$ that indicate the presence of a vortex at location $\bm{r}=m\,\bm{A}_{1}+n\,\bm{A}_{2}$ for $(m,n)\in\mathbb{Z}^{2}$. For dual configurations [see Fig. 1 (b) and (c)], we define the vectors $\bm{A}_{1}$ and $\bm{A}_{2}$ to indicate the absence of a vortex at position $\bm{r}$. In the following we shall refer to the parallelogram built from the vectors $\bm{A}_{1}$ and $\bm{A}_{2}$ as the geometric unit cell (GUC). It is not necessarily the same as the Hamiltonian unit cell (HUC), spanned by the Bravais lattice vectors $\bm{a}_{1}$ and $\bm{a}_{2}$, and which depends on the gauge choice for the $u_{jk}$ variables.
@@
Table 2: Four families built from translation ($P_{w}=\pm 1$) and inversion ($\mathcal{P}^{2}=\pm\mathbbm{1}$) symmetries. The geometric (GUC) and Hamiltonian (HUC) unit cells are indicated. The direct and dual configurations considered in this article occupy three families.
@@
Throughout the rest of this article, unless otherwise stated, we shall define our energy units such that $J=1$, with the assumption that $J$ is nonzero and positive.
@@ III Translation and inversion symmetries
Whether or not the gauge choice for the $u_{jk}$ variables is invariant under translations of $\bm{A}_{1}$ and $\bm{A}_{2}$ is found to depend on the parity of the vortex number $P_{w}$ inside the GUC [15, 14]. When $P_{w}=+1$, there is an even number of vortices in the GUC. In this case, there exists a gauge choice invariant under translations of $\bm{A}_{1}$ and $\bm{A}_{2}$ and the HUC is the same as the GUC [8, 15]. By contrast, when $P_{w}=-1$, there is an odd number of vortices in the GUC, and the HUC must be double that of the GUC. Importantly, it has been shown that when $P_{w}=-1$, the system can only exhibit even-valued Chern numbers, while if $P_{w}=+1$ both even and odd Chern numbers are possible. The direct vortex configurations always exhibit $P_{w}=-1$, while the dual configurations can either exhibit $P_{w}=+1$ or $P_{w}=-1$ [15]. A vortex configuration with $P_{w}=+1$ is shown in Fig. 1(b), while vortex configurations with $P_{w}=-1$ can be seen in Figs. 1(a) and (c).
@@
In Table 2, we classify different families based on the signs of $P_{w}=\pm 1$ and $\mathcal{P}^{2}=\pm\mathbbm{1}$. There are neither direct nor dual configurations for which $P_{w}=+1$ and $\mathcal{P}^{2}=+\mathbbm{1}$. It was implicit in the analysis of Ref. [15] that for the vortex configurations studied there, the gauge transformation can be chosen to have the periodicity of the Hamiltonian (i.e. that the operator $\mathcal{P}$ commutes with the translation operators defining the HUC). However, this might not be true for other vortex configurations on the honeycomb lattice or for the Kitaev model on other lattices, such as the square-octagon [23] or the triangle-dodecagon [26, 27, 24, 28].
@@ IV Effective low-energy models
Figure 3: Energy versus IDOS of $H$ plotted in blue for: (a) direct ($\rho=1/21$), (b) dual even ($\rho=15/16$), and (c) dual odd ($\rho=20/21$), all with $\kappa=1$ (energy units are such that $J=1$). In (a), the thick red line gives the IDOS of the vortex-free background, while for (b) and (c) it gives that of the vortex-full background. The dotted grey lines indicate the total bandwidth $\delta_{\rm v}$ of the vortex bands, while the dashed grey lines indicate the energy gap $\Delta_{\rm b}$ of the background bands. The second row shows a zoom on the vortex bands (there are two for the direct and dual odd configurations, and four for the dual even configurations).
@@
Previous works have shown that sufficiently sparse direct vortex configurations can be described by an effective Majorana tight-binding model that accurately reproduces their spectra and Chern numbers [12, 13]. The intuition behind this model follows from analyzing the behavior of isolated vortices: given a vortex-free background, exchanging $u_{jk}\to-u_{jk}$ for a single link variable introduces two adjacent vortices into the system. When these vortices are sufficiently separated from each other, these $w_{p}=-1$ plaquettes each host a single unpaired MZM [5]. The spectrum of a dilute direct vortex lattice against a vortex-free background can then be understood as arising from a set of high-energy bands coming from the background (“background bands”) and a set of low-energy bands coming from the tunneling of Majorana fermions between these zero modes (“vortex bands”). This can be seen, for example, in Fig. 3(a) where we plot the IDOS of the Hamiltonian in Eq. (1) exhibiting a direct vortex configuration with density $\rho=1/21$. The total Chern number of the system in this case can then be decomposed as $\nu=\nu_{\rm b}+\nu_{\rm v}$, where $\nu_{\rm b}=1$ and $\nu_{\rm v}$ are the Chern numbers for the vortex-free background bands and the vortex bands respectively.
@@
In this section we expand upon the above analysis and derive low-energy effective models for the two families of dual vortex configurations. These cases are distinct from the direct configurations as an isolated $w_{p}=+1$ plaquette within a vortex-full background does not host an unpaired MZM. Instead, in the dual case, the even Chern number of the vortex-full background implies that an isolated “dual vortex” (i.e. a vortex-free plaquette surrounded by vortex-full plaquettes) will host a pair of coupled Majorana fermions equivalent to a single complex fermion mode with finite energy. The tunneling between these isolated finite-energy modes in the dilute limit then gives rise to a set of low-energy bands separated from the vortex-full background (we will still refer to these bands as “vortex bands”). This can be seen in Figs. 3(b) and (c), where we plot the IDOS for $\rho=15/16$ (dual even) and $\rho=20/21$ (dual odd). The number of low-energy bands (either two or four) depends on the symmetry family, as discussed below.
@@ IV.1 Warm-up: direct case
where $\gamma_{i}$ are Majorana operators defined on the $w_{p}=-1$ plaquettes of the honeycomb lattice and $(t_{1},t_{2})\in\mathbb{R}^{2}$ represent the hopping amplitudes between nearest and next-nearest neighbor vortices, respectively. As explained in Ref. [12], one needs these two terms to reproduce the Chern numbers observed. The $\mathbb{Z}_{2}$ gauge variables $s_{jk}=\pm 1$ are chosen such that the total flux accumulated by a Majorana fermion hopping counter-clockwise on a closed triangular circuit of $t_{1}$ links is $\phi_{1}=\pi/2$, the flux enclosed in a closed circuit of $t_{2}$ links is $\phi_{2}=-\pi/2$, and the flux enclosed in a closed circuit formed by both $t_{1}$ and $t_{2}$ links is $\phi_{1,2}=\pi/2$ (see Fig 4). In contrast with Ref. [12], we choose a periodic pattern of $s_{jk}$ variables that both satisfies these requirements and gives rise to only two sites per unit cell. Utilizing the gauge choice defined in Fig. 4, we Fourier transform the Hamiltonian in Eq. (3) to obtain
Figure 4: Direct: Unit cell of the effective model defined in Eq. (3). Grey and hatched hexagons indicate the two inequivalent sites in the unit cell of the effective model. The directions of the blue dashed and red dotted arrows represent the sign of the gauge variables for the nearest ($t_{1}$) and next-nearest ($t_{2}$) neighbor hopping terms, respectively: If an arrow points from vortex site $k$ to vortex site $j$, then $s_{jk}=-s_{kj}=+1$.
Figure 5: Direct: (a) Phase diagram and Chern numbers $\nu_{\rm v}$ of the effective model given in Eq. (3), adapted from Ref. [12]. The phase boundaries are $t_{2}=t_{1}$, $-t_{1}$, and $-t_{1}/2$. (b) Comparison of the IDOS of the effective model and that of the vortex bands of $H$ (both IDOS are normalized between 0 to 1). Here we have $\delta_{\rm v}/\Delta_{\rm b}\simeq 0.009$. The parameters are $t_{1}=-0.0028$, and $t_{2}=-0.0002$. These parameter values are indicated by the red dot in the phase diagram of (a), implying that $\nu_{\rm v}=+1$, in agreement with the microscopic model.
@@
Here we have defined the nearest neighbor directions as $\bm{a}_{1}/2$, $\bm{a}_{2}$, and $\bm{a}_{3}=\bm{a}_{2}-\bm{a}_{1}/2$, as well as the next-nearest neighbor directions as $\bm{b}_{1}=\bm{a}_{1}-\bm{a}_{2}$, $\bm{b}_{2}=\bm{a}_{2}+\bm{a}_{3}$, and $\bm{b}_{3}=\bm{a}_{2}+\bm{a}_{1}/2$. The above Majorana Hamiltonian in Eq. (11) can be transformed into a Bogoliubov-de Gennes Hamiltonian by a $\bm{k}$-independent unitary transformation, revealing that the system is equivalent to an anisotropic $p_{x}+ip_{y}$ topological superconductor (see e.g. Sec. 16.3 in Ref. [29]).
In Fig. 5(a) we reproduce the phase diagram of this model from Ref. [12], which features the possible vortex Chern numbers $\nu_{\rm v}=\pm 1,\pm 3$. We call $E_{\rm min}(t_{1},t_{2})$ and $E_{\rm max}(t_{1},t_{2})$ the minimum and maximum positive energy eigenvalues of Eq. (11), which are equal to
@@
Here, we used these extrema to optimize the fitting parameters $t_{1}$ and $t_{2}$ by imposing the proper bound of the IDOS (they will also be used in the following section). To give an example of the accuracy of this model, in Fig. 5(b) we plot the IDOS of the low-energy bands of the Kitaev honeycomb model in Eq. (1) along with that of the effective model given in Eq. (4). Here, we fit the parameters $t_{1}$ and $t_{2}$ by matching the IDOS (and Chern number) corresponding to Eq. (4) with that of Eq. (1). It is possible that multiple choices of $t_{1}$ and $t_{2}$ may result in the same IDOS as that of Eq. (1). In this case, as argued in Ref. [12], the relative magnitudes of these parameters may be predicted by extracting $t_{1}$ and $t_{2}$ from the energy splitting of MZMs obtained by a numerical calculation with two spatially-separated vortices in a vortex-free background [12]. While this method is less precise than exactly matching the IDOS, it has the advantage that it provides a unique solution. Therefore, both approaches should be combined.

### 05 Mar 2024 v2

@@ I Introduction
The characterization of free fermion phases via topological invariants represents a major achievement in the understanding of quantum matter [1, 2, 3, 4]. Given the spatial dimension, these invariants may be determined through an analysis of global symmetries such as time-reversal (TRS), particle-hole, and chiral symmetries, in a classification scheme known as the tenfold way. In 2006, Kitaev introduced an exactly solvable spin-1/2 model on a honeycomb lattice that can be mapped to a free Majorana-fermion problem in a static $\mathbb{Z}_{2}$ gauge field [5]. The gauge field results in the possibility of static $\pi$ fluxes, called vortices, in each plaquette and the problem separates into independent vortex sectors. In the isotropic limit and in the vortex-free sector where the ground state lies (see below for details), the spectrum is gapless, but acquires a gap and a nonzero Chern number $\nu$ when TRS is broken [5]. To break TRS while preserving the integrability of the model, Kitaev introduced a three-spin term that results in a next-nearest neighbor hopping of the Majorana fermions. The corresponding model, closely related to Haldane’s model [6], has $\nu=\pm 1$ and has been the subject of many studies (see, e.g., Refs. [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]).
@@
The “duals” to these triangular vortex configurations correspond to dense vortex configurations, with density $\rho=(n-1)/n$, where the plaquettes that do not contain vortices form a triangular pattern against a background of plaquettes that all contain vortices. These dense vortex configurations have recently been studied numerically in Refs. [16, 17], and can be divided into two families according to their behavior with respect to translation and inversion symmetry. We shall refer to these two families as “dual even” (if $n$ is even such as $\rho=3/4,11/12,15/16,...$) and “dual odd” (if $n$ is odd such as $\rho=2/3,8/9,11/12,...$) in the following [see Figs. 1(b) and 1(c)] [8, 17]. The dual even family exhibits gapped phases with even Chern numbers, separated by extended gapless phases (spin metals with a Majorana-Fermi surface [26, 27]), whereas the dual odd family has gapped phases with even or odd Chern numbers, separated by gapless points. These dual configurations are not describable by the effective model introduced in Refs. [12, 13].
In this paper, we construct two different effective models describing the dual even and dual odd families. These models describe the hopping of pairs of coupled Majorana fermions on a triangular lattice, where the lattice sites correspond to the absences of vortices within a vortex sea. These effective models allow us to understand the variety of Chern numbers and the nature of gapless phases observed. The paper is organized as follows: in Sec. II we briefly introduce the Kitaev honeycomb model in the Majorana fermion language and discuss the behaviors of the vortex-free and vortex-full sectors. In Sec. III, we identify three distinct symmetry families for direct and dual vortex configurations based on translation and inversion symmetries. In Sec. IV, we show that when the low-energy bands of the Kitaev honeycomb model are well-separated from the high-energy bands, they may be accurately described by an effective tight-binding model for all three of these families. We first review the construction of the low-energy model discussed in Refs. [12, 13] for the direct case, and then introduce new effective models for the dual even and dual odd cases. To demonstrate the accuracy of these models, we directly compare their integrated density of states (IDOS) with that of the original Kitaev honeycomb model. Finally, in Sec. V, we discuss our results, conclude and give perspectives. Appendices provide details on specific points such as the validity of the effective models (Appendix A) and some possible dispersion relations of vortex bands (Appendix B).
@@ II Kitaev honeycomb model
Here the $u_{jk}$ parameters are $\mathbb{Z}_{2}$ gauge variables defined such that $u_{jk}=-u_{kj}=\pm 1$ for each link of the lattice. We assume here that the coupling $J$ is the same in all three directions, corresponding to the isotropic limit of the original spin model. This Hamiltonian describes noninteracting Majorana fermions coupled to a static $\mathbb{Z}_{2}$ gauge field. The term $J$ describes the nearest-neighbor hopping, whereas $\kappa$ denotes the strength of the three-spin term that breaks TRS, corresponding to next-nearest-neighbor hopping (see Fig. 2). In the following, without loss of generality, we will assume that $J$ and $\kappa$ are nonnegative.
For each plaquette $p$ of the honeycomb lattice, we define the $\mathbb{Z}_{2}$ variable $w_{p}=\prod_{(j,k)\in p}u_{jk}={\rm e}^{i\phi_{p}}$, where $\phi_{p}$ is the flux threading the plaquette. Here, for each plaquette, the pair of sites $(j,k)$ is taken such that $k$ always belongs to one sublattice of the honeycomb lattice, and $j$ belongs to the other. A value of $w_{p}=-1$ ($+1$) indicates a flux $\phi_{p}=\pi$ (00) in the plaquette $p$, meaning the presence (absence) of a vortex. Two sets of $\mathbb{Z}_{2}$ link variables $u_{jk}$ defined on the lattice are said to be equivalent if they correspond to the same configuration of plaquette variables, or “vortex configuration.” In analyzing Eq. (1), one can then restrict to a particular vortex sector labeled by the pattern of $w_{p}$’s. In Fig. 1, we show several examples of vortex configurations, where the $w_{p}=-1$ plaquettes are shaded grey.
Direct vortex configurations [see Fig. 1 (a)] are defined by two vectors $\bm{A}_{1}$ and $\bm{A}_{2}$ that indicate the presence of a vortex at location $\bm{r}=m\,\bm{A}_{1}+n\,\bm{A}_{2}$ for $(m,n)\in\mathbb{Z}^{2}$. For dual configurations [see Figs. 1(b) and 1(c)], we define the vectors $\bm{A}_{1}$ and $\bm{A}_{2}$ to indicate the absence of a vortex at position $\bm{r}$. In the following we shall refer to the parallelogram built from the vectors $\bm{A}_{1}$ and $\bm{A}_{2}$ as the geometric unit cell (GUC). It is not necessarily the same as the Hamiltonian unit cell (HUC), spanned by the Bravais lattice vectors $\bm{a}_{1}$ and $\bm{a}_{2}$, and which depends on the gauge choice for the $u_{jk}$ variables.
@@
Table 2: Four families built from translation ($P_{w}=\pm 1$) and inversion ($\mathcal{P}^{2}=\pm\mathbbm{1}$) symmetries. The geometric (GUC) and Hamiltonian (HUC) unit cells are indicated. The direct and dual configurations considered in this paper occupy three families.
@@
Throughout the rest of this paper, unless otherwise stated, we shall define our energy units such that $J=1$, with the assumption that $J$ is nonzero and positive.
@@ III Translation and inversion symmetries
Whether or not the gauge choice for the $u_{jk}$ variables is invariant under translations of $\bm{A}_{1}$ and $\bm{A}_{2}$ is found to depend on the parity of the vortex number $P_{w}$ inside the GUC [17, 16]. When $P_{w}=+1$, there is an even number of vortices in the GUC. In this case, there exists a gauge choice invariant under translations of $\bm{A}_{1}$ and $\bm{A}_{2}$ and the HUC is the same as the GUC [8, 17]. By contrast, when $P_{w}=-1$, there is an odd number of vortices in the GUC, and the HUC must be double that of the GUC. Importantly, it has been shown that when $P_{w}=-1$, the system can only exhibit even-valued Chern numbers, while if $P_{w}=+1$ both even and odd Chern numbers are possible. The direct vortex configurations always exhibit $P_{w}=-1$, while the dual configurations can either exhibit $P_{w}=+1$ or $P_{w}=-1$ [17]. A vortex configuration with $P_{w}=+1$ is shown in Fig. 1(b), while vortex configurations with $P_{w}=-1$ can be seen in Figs. 1(a) and 1(c).
@@
In Table 2, we classify different families based on the signs of $P_{w}=\pm 1$ and $\mathcal{P}^{2}=\pm\mathbbm{1}$. There are neither direct nor dual configurations for which $P_{w}=+1$ and $\mathcal{P}^{2}=+\mathbbm{1}$. It was implicit in the analysis of Ref. [17] that for the vortex configurations studied there, the gauge transformation can be chosen to have the periodicity of the Hamiltonian (i.e., that the operator $\mathcal{P}$ commutes with the translation operators defining the HUC). However, this might not be true for other vortex configurations on the honeycomb lattice or for the Kitaev model on other lattices, such as the square-octagon [26] or the triangle-dodecagon [29, 30, 27, 31].
@@ IV Effective low-energy models
Figure 3: Energy vs IDOS of $H$ plotted in blue for: (a) direct ($\rho=1/21$), (b) dual even ($\rho=15/16$), and (c) dual odd ($\rho=20/21$), all with $\kappa=1$ (energy units are such that $J=1$). In (a), the thick red line gives the IDOS of the vortex-free background, while for (b) and (c) it gives that of the vortex-full background. The dotted grey lines indicate the total bandwidth $\delta_{\rm v}$ of the vortex bands, while the dashed grey lines indicate the energy gap $\Delta_{\rm b}$ of the background bands. The second row shows a zoom on the vortex bands (there are two for the direct and dual odd configurations, and four for the dual even configurations).
@@
Previous works have shown that sufficiently sparse direct vortex configurations can be described by an effective Majorana tight-binding model that accurately reproduces their spectra and Chern numbers [12, 13]. The intuition behind this model follows from analyzing the behavior of isolated vortices: given a vortex-free background, exchanging $u_{jk}\to-u_{jk}$ for a single link variable introduces two adjacent vortices into the system. When these vortices are sufficiently separated from each other, these $w_{p}=-1$ plaquettes each host a single unpaired MZM [5]. The spectrum of a dilute direct vortex lattice against a vortex-free background can then be understood as arising from a set of high-energy bands coming from the background (“background bands”) and a set of low-energy bands coming from the tunneling of Majorana fermions between these zero modes (“vortex bands”). This can be seen, for example, in Fig. 3(a) where we plot the IDOS of the Hamiltonian in Eq. (1) exhibiting a direct vortex configuration with density $\rho=1/21$. The IDOS $N(E)$ is defined as the number of states below a given energy $E$. Here, for practical use, we rather plot $E(N)$. The total Chern number of the system in this case can then be decomposed as $\nu=\nu_{\rm b}+\nu_{\rm v}$, where $\nu_{\rm b}=1$ and $\nu_{\rm v}$ are the Chern numbers for the vortex-free background bands and the vortex bands, respectively.
@@
In this section, we expand upon the above analysis and derive low-energy effective models for the two families of dual vortex configurations. These cases are distinct from the direct configurations as an isolated $w_{p}=+1$ plaquette within a vortex-full background does not host an unpaired MZM. Instead, in the dual case, the even Chern number of the vortex-full background implies that an isolated “dual vortex” (i.e., a vortex-free plaquette surrounded by vortex-full plaquettes) will host a pair of coupled Majorana fermions equivalent to a single complex fermion mode with finite energy. The tunneling between these isolated finite-energy modes in the dilute limit then gives rise to a set of low-energy bands separated from the vortex-full background (we will still refer to these bands as “vortex bands”). This can be seen in Figs. 3(b) and 3(c), where we plot the IDOS for $\rho=15/16$ (dual even) and $\rho=20/21$ (dual odd). The number of low-energy bands (either two or four) depends on the symmetry family, as discussed below.
@@ IV.1 Warm-up: direct case
where $\gamma_{i}$ are Majorana operators defined on the $w_{p}=-1$ plaquettes of the honeycomb lattice and $(t_{1},t_{2})\in\mathbb{R}^{2}$ represent the hopping amplitudes between nearest and next-nearest-neighbor vortices, respectively. As explained in Ref. [12], one needs these two terms to reproduce the Chern numbers observed. The $\mathbb{Z}_{2}$ gauge variables $s_{jk}=\pm 1$ are chosen such that the total flux accumulated by a Majorana fermion hopping counter-clockwise on a closed triangular circuit of $t_{1}$ links is $\phi_{1}=\pi/2$, the flux enclosed in a closed circuit of $t_{2}$ links is $\phi_{2}=-\pi/2$, and the flux enclosed in a closed circuit formed by both $t_{1}$ and $t_{2}$ links is $\phi_{1,2}=\pi/2$ (see Fig 4). In contrast with Ref. [12], we choose a periodic pattern of $s_{jk}$ variables that both satisfies these requirements and gives rise to only two sites per unit cell. Utilizing the gauge choice defined in Fig. 4, we Fourier transform the Hamiltonian in Eq. (3) to obtain
Figure 4: Direct. Unit cell of the effective model defined in Eq. (3). Grey and hatched hexagons indicate the two inequivalent sites in the unit cell of the effective model. The directions of the blue dashed and red dotted arrows represent the sign of the gauge variables for the nearest ($t_{1}$) and next-nearest ($t_{2}$) neighbor hopping terms, respectively: if an arrow points from vortex site $k$ to vortex site $j$, then $s_{jk}=-s_{kj}=+1$.
Figure 5: Direct. (a) Phase diagram and Chern numbers $\nu_{\rm v}$ of the effective model given in Eq. (3), adapted from Ref. [12]. The phase boundaries are $t_{2}=t_{1}$, $-t_{1}$, and $-t_{1}/2$. (b) Comparison of the IDOS of the effective model and that of the vortex bands of $H$ (both IDOS are normalized between 0 to 1). Here we have $\delta_{\rm v}/\Delta_{\rm b}\simeq 0.009$. The parameters are $t_{1}=-0.0028$, and $t_{2}=-0.0002$. These parameter values are indicated by the red dot in the phase diagram of (a), implying that $\nu_{\rm v}=+1$, in agreement with the microscopic model.
@@
Here we have defined the nearest-neighbor directions as $\bm{a}_{1}/2$, $\bm{a}_{2}$, and $\bm{a}_{3}=\bm{a}_{2}-\bm{a}_{1}/2$, as well as the next-nearest neighbor directions as $\bm{b}_{1}=\bm{a}_{1}-\bm{a}_{2}$, $\bm{b}_{2}=\bm{a}_{2}+\bm{a}_{3}$, and $\bm{b}_{3}=\bm{a}_{2}+\bm{a}_{1}/2$. The above Majorana Hamiltonian in Eq. (11) can be transformed into a Bogoliubov-de Gennes Hamiltonian by a $\bm{k}$-independent unitary transformation, revealing that the system is equivalent to an anisotropic $p_{x}+ip_{y}$ topological superconductor (see, e.g., Sec. 16.3 in Ref. [32]).
In Fig. 5(a), we reproduce the phase diagram of this model from Ref. [12], which features the possible vortex Chern numbers $\nu_{\rm v}=\pm 1,\pm 3$. We call $E_{\rm min}(t_{1},t_{2})$ and $E_{\rm max}(t_{1},t_{2})$ the minimum and maximum positive energy eigenvalues of Eq. (11), which are equal to
@@
Here, we used these extrema to optimize the fitting parameters $t_{1}$ and $t_{2}$ by imposing the proper bound of the IDOS (they will also be used in the following section). To give an example of the accuracy of this model, in Fig. 5(b) we plot the IDOS of the low-energy bands of the Kitaev honeycomb model in Eq. (1) along with that of the effective model given in Eq. (4). Here, we fit the parameters $t_{1}$ and $t_{2}$ by matching the IDOS (and Chern number) corresponding to Eq. (4) with that of Eq. (1). It is possible that multiple choices of $t_{1}$ and $t_{2}$ may result in the same IDOS as that of Eq. (1). In this case, as argued in Ref. [12], the relative magnitudes of these parameters may be predicted by extracting $t_{1}$ and $t_{2}$ from the energy splitting of MZMs obtained by a numerical calculation with two spatially separated vortices in a vortex-free background [12]. While this method is less precise than exactly matching the IDOS, it has the advantage that it provides a unique solution. Therefore both approaches should be combined.